OpenStudy (watchmath):

Interesting problem anyone? Compute \[\lim_{n\to\infty}\frac{1\cdot 1!+2\cdot 2!+\cdots+n\cdot n!}{(n+1)!}\]

OpenStudy (anonymous):

guess or proof?

OpenStudy (anonymous):

just kidding.

OpenStudy (watchmath):

guess is welcome too :D

OpenStudy (anonymous):

1

OpenStudy (watchmath):

Good guess!! :D

OpenStudy (anonymous):

think i even know what it is, but the proof is eluding me so now i have something to think about for the afternoon.

OpenStudy (anonymous):

closed form i mean.

OpenStudy (anonymous):

won't post until i prove it.

OpenStudy (watchmath):

alrighty!

OpenStudy (anonymous):

Can we not just factor out an n! from each term in the top, and rewrite it as: \[\lim_{n \rightarrow \infty} \frac{n\cdot n![\frac{1}{n\cdot n!} + \frac{2 \cdot 2!}{n\cdot n!} + ... + 1]}{n\cdot n!(1+\frac{1}{n})} = \lim_{n \rightarrow \infty} \frac{\frac{1}{n\cdot n!} + \frac{2 \cdot 2!}{n\cdot n!} + ... + 1}{(1+\frac{1}{n})} = \frac{1}{1} = 1\]

OpenStudy (anonymous):

I guess it's not clear that the sum of all those n terms will go to 0 faster than the number of terms are growing but it seems reasonable to me.

OpenStudy (watchmath):

Yes, I think it works. We just need to show more rigorously that \[\lim_{n\to\infty}\frac{(n-r)\cdot(n-r)!}{n\cdot n!}=0\] for all \(0<r<n\).

OpenStudy (anonymous):

yeah, I'm looking at that now actually ;)

OpenStudy (anonymous):

actually just showing it works for r=1 will show it for the rest

OpenStudy (watchmath):

Agree :).

OpenStudy (anonymous):

I found it

OpenStudy (anonymous):

\[(n-1)! = \frac{n!}{n} \rightarrow \lim_{n\to\infty}\frac{(n-1)\cdot (n-1)!}{n\cdot n!} = \lim_{n\to\infty}\frac{(n-1)\cdot n!}{n^2 n!} = \lim_{n\to\infty}\frac{(n-1)}{n^2} = 0\]

OpenStudy (anonymous):

i think it is just \[\frac{(n+1)!-1}{(n+1)!}\]yes?

OpenStudy (watchmath):

Awesome satellite! :D

OpenStudy (anonymous):

proof by induction as soon as i figure it out.

OpenStudy (anonymous):

My method doesn't work?

OpenStudy (watchmath):

Your method works! Satellite shows that the expression is in fact equal to \(\frac{(n+1)!-1}{(n+1)!}\)

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

well actually satellite has not shown anything. i just said it.

OpenStudy (anonymous):

so no credit yet that is for sure.

OpenStudy (anonymous):

Fun problem though

OpenStudy (anonymous):

i thought it would be simple induction. but i am getting stuck because i keep getting \[n(n+1)!+2(n+1)!\] and i need this to be \[(n+2)!\] which, if it is, is not clear to me.

OpenStudy (anonymous):

any hints?

OpenStudy (watchmath):

Are you sure you want a hint. I am afraid it will soil the fun :).

OpenStudy (anonymous):

Just factor

OpenStudy (anonymous):

ok no hint. perhaps induction is not way to go?

OpenStudy (anonymous):

oh lord i get the dumb guy award

OpenStudy (watchmath):

ok, half hint. The summation on the top can be made into a telescoping sum.

OpenStudy (anonymous):

hehe.. sometimes we can't see the forest for the trees. Happens to us all.

OpenStudy (anonymous):

polpak got it. proof by induction done.

OpenStudy (anonymous):

but now i have to think about the telescope.

OpenStudy (watchmath):

\((n+1)!-1=(n+1)!-1!\) :D

OpenStudy (watchmath):

I think a hint won't hurt now :D. \(n\cdot n!=((n+1)-1)\cdot n!\)